An MTN cellular-phone user notices that her cellular phone accounts for three
consecutive months were significantly different. In May, her account was one and a half
times more than her account in June. Her account in July was R 50 more than her account in
June. In total, she spent R575 in cellular-phone accounts in the three months. What was her
account for each month?
1
Expert's answer
2019-04-26T14:12:46-0400

"x_1" - R was spent in May

"x_2" - R was spent in June

"x_3" - R was spent in July


Construct the system of equations:



"x_1 = 1.5 * x_2""x_2 = x_3 - 50""x_1 + x_2 + x_3 = 575"



Since "x_2 = x_3 - 50" and "x_1 = 1,5 * x_2" thus:


"1.5 * (x_3 - 50) + x_3 - 50 + x_3 = 575"

"1.5 * x_3 + x_3 + x_3 = 575 + 50 + 75"

"3.5 * x_3 = 700"

"x_3 = 200"


Find x2 :

"x_2 = x_3 - 50 = 200 - 50 = 150"


Find x1 :

"x_1 = 1.5 * x_2 = 1.5 * 150 = 225"



Answer:


May: x1 = 225

June: x2 = 150

July: x3 = 200